### v1

# Learning in Strategic Queuing Systems with Small Buffers

Abstract
Routers in networking use simple learning algorithms to find the best way to deliver packets to their desired destination. This simple, myopic and distributed decision system makes large queuing systems simple to operate, but at the same time, the system needs more capacity than would be required if all traffic were centrally coordinated. In a recent paper, Gaitonde and Tardos (EC 2020 and JACM 2023) initiate the study of such systems, modeling them as an infinitely repeated game in which routers compete for servers and the system maintains a state (number of packets held by each queue) that results from outcomes of previous rounds. Queues get to send a packet at each step to one of the servers, and servers attempt to process only one of the arriving packets, modeling routers. However, their model assumes that servers have no buffers at all, so queues have to resend all packets that were not served successfully. They show that, even with hugely increased server capacity relative to what is needed in the centrally-coordinated case, ensuring that the system is stable requires the use of timestamps and priority for older packets.
We consider a system with two important changes, which make the model more realistic: first we add a very small buffer to each server, allowing the server to hold on to a single packet to be served later (even if it fails to serve it); and second, we do not require timestamps or priority for older packets. Our main result is to show that when queues are learning, a small constant factor increase in server capacity, compared to what would be needed if centrally coordinating, suffices to keep the system stable, even if servers select randomly among packets arriving simultaneously.
This work contributes to the growing literature on the impact of selfish learning in systems with carryover effects between rounds: when outcomes in the present round affect the game in the future.
1 Introduction
In this paper, we model routers in networking as agents using simple learning algorithms to find the best way to get their packets served. This simple, myopic and distributed multi-agent decision system makes large queuing systems simple to operate, but at the same time, the system needs more capacity than would be required if all traffic were centrally coordinated. In a recent paper, Gaitonde and Tardos (2020, 2023), the authors initiated the study of such systems, modeling them as an infinitely repeated game in which routers compete for servers and the system maintains a state (number of packets held by each queue) that results from outcomes of previous rounds. Routers get to send a packet at each step at which they received a packet or have one in their queue [^1] The players are routers with states, the length of their queues; the terms router and queue are used interchangeably. to one of the servers, and each server attempts to process only one of the packets arriving to it. However, the model of Gaitonde and Tardos (2020, 2023) assumes that servers have no buffers at all, so queues have to resend all packets that were not served successfully. They show that, in their system, even with hugely increased server capacity relative to what is needed in the centrally-coordinated case, the use of timestamps and priority for older packets is required to ensure that the system is stable, i.e., that queue lengths do not diverge with time (see Section 2 for a formal definition).
In networking systems, servers accepting packets typically have a very small buffer and can hold on to a few packets to be served later. In this paper, we consider the analog of the model of Gaitonde and Tardos (2020, 2023) in such systems with a tiny buffer at each server. We show that even with a buffer capable of holding only a single packet, the service capacity of the system greatly increases even when agents are learning. To see this, consider a system with a single queue receiving packets at a high rate, say with probability $\lambda=\frac{1}{2}$ at each time step, and multiple servers, each with a service rate (the probability of successfully serving a packet), say $\mu=\frac{1}{3}$. Without buffers, the condition for stability of such a system is that there is a single server with a service rate of $\mu>\lambda$, as a queue can only send a packet to one server at a time. So even with a large number of servers, each with service rate $\mu=\frac{1}{3}$, and thus, an arbitrarily large sum of service capacities, it would not be feasible to keep the system stable. Also, the learning algorithm of this single queue is facing a classical multi-arm bandit problem, aiming to learn which of the servers is the best one to send their packets to. By contrast, we observe that with a buffer of even just one packet at each server, a large enough number of servers with low capacity will always ensure stability, as the learning agent can now take advantage of low capacity servers as well, sending packets to them rarely enough such that they are likely to clear in time before the next packet is sent to them.
Back to our example of servers with $\mu=\frac{1}{3}$ and $\lambda=\frac{1}{2}$, which, as we saw above, is not stable without buffers for any number of servers, with buffers, two such servers will make the system stable if the queue alternates sending its packets to them. In light of this observation, the goal of a learning algorithm even with just one queue and many servers is now more complex: it should learn to distribute its packets in a way that takes advantage of the capacity of all the servers.
Our main result, formally stated in Theorem 1, is to show that when multiple queues compete for service while using learning algorithms to identify good servers, a small constant-factor increase in server capacity—relative to what would be required under centralized coordination—suffices to maintain the system stable, even if servers randomly select among simultaneously arriving packets. Specifically, suppose we have $n$ queues with packet arrival rates of $\lambda_{1},\ldots,\lambda_{n}$, and $m$ servers of service rates $\mu_{1},\ldots,\mu_{m}$. Obviously, we must have $\sum_{i}\lambda_{i}<\sum_{j}\mu_{j}$, else the system cannot be stable even if fully coordinated. We show that this condition is not strong enough to ensure stability even with full coordination due to the small buffers. [^2] This condition allows to create a fractional assignment such that $\sum_{j}x_{ij}>\lambda_{i}$ and $\sum_{i}x_{ij}<\mu_{j}$. Using the matching decomposition of this assignment in a coordinated schedule will guarantee a bounded expected number of packets both at the queues and servers. However, our result shows that if $\lambda_{i}<\frac{1}{2}$ for all $i$, all queues use a learning algorithm that guarantees the no-regret condition with high probability, and $3\sum_{i}\lambda_{i}<\sum_{j}\mu_{j}$, this guarantees that the system remains stable. While we do not know whether the factor of 3 in the required capacity constraint is tight, we show that to guarantee that a no-regret outcome of learning by the queues remains stable, it is required to have at least $2\sum_{i}\lambda_{i}<\sum_{j}\mu_{j}$ (see Remark 1 and Theorem 3 for more details).
Related Work.
The classical focus of work on scheduling in queuing systems is aimed at finding schedules that achieve optimal throughput (see, e.g., the textbook of Shortle et al. (2018)). For work evaluating efficiency loss due to selfishness in different classical queuing systems, see the book of Hassin Hassin (2020) and survey of Hassin and Haviv (2003). Closer to our motivation, there is a growing literature that aims to understand how systems perform when the queues use simple learning algorithm to find good service. Krishnasamy et al. (2016) considers a queue using a no-regret learning algorithm to find what may be the best servers, but does not consider multiple queues competing for service. Their primary goal is to study the expected size of the queue of packets as a function of time. They also extend the result to the case of multiple queues scheduled by a single coordinated scheduling algorithm, assuming there is a perfect matching between queues and optimal servers that can serve them. Sentenac et al. (2021) and Freund et al. (2022) extended this work to decentralized learning dynamics in bipartite queuing systems that attain near-optimal performance without the matching assumption. For a survey on the role of learning and information in queuing systems see Walton and Xu (2021).
In our work, we focus on a game-theoretic model: queues using learning algorithms to best distribute their packets to get good service, while also competing for servers. We assume that each queue separately learns to selfishly make sure its own packets are served at the highest possible rate, offering a strategic model of scheduling packets in a queuing system. Closest to our model from this literature is the work Gaitonde and Tardos (2020, 2023), who consider the same bipartite model of queues and servers as we do, but without buffers. They show the exact condition to make such a system stable with central coordination of packets and prove that no-regret learning by the queues guarantees stability of the system if it has double the capacity needed for central coordination, assuming packets carry a timestamp and servers choose the oldest packet to serve. Fu et al. (2022) extended this work to a general network. Baudin et al. (2023) propose an alternative, episodic queuing system where agents have incentives to hold jobs in an episode before sending to a central server, but suffer penalties should their jobs not be completed before the end of the episode. They show that both equilibrium and no-regret outcomes ensure stability, so long as these costs are sufficiently large.
In the works discussed so far, the servers receiving the packets have no buffers: all unserved packets are held in a queue and get resent to be served later. In queuing systems aimed at modeling networking, the receivers (servers) each do have a very small buffer, and can hold on to a couple of packets to be served later. It turns out that even having a buffer for a single packet significantly changes the effective service capacity of the network, as already explained above.
In the case without such buffers, it was feasible to exactly characterize the capacity needed to be able to make the system stable with central coordination by an elegant use of linear programming duality (see Fu et al. (2022)). In the presence of limited buffer capacity at the servers, the rate at which a server accepts a packet depends on the state of the buffers. This is in contrast to a system without buffers, where the probability that a server accepts a packet is the service rate of the server, and is stable over time. There is some recent work considering queuing systems with such evolving service probabilities, see for example Grosof et al. (2024b, a). The goal of that work is to characterize the exact condition that allows the system to remain stable with central coordination of the schedule. Our work is the first to consider stability in such a system assuming each queue independently aims to optimize its own service rate running a no-regret learning algorithm.
There is a large literature on bounding the price of anarchy for various games, beginning with Koutsoupias and Papadimitriou (1999). These results extend to analyzing outcomes in repeated games assuming players are no-regret learners Balcan et al. (2013); Roughgarden (2015); Syrgkanis and Tardos (2013). However, these works make the strong assumption that games in different rounds are independent. By contrast, different rounds in queuing games are not independent: the number of packets in the system depends on the success in previous periods. Work on games with carryover effects, or context, are considered in multi-agent reinforcement learning as well as in Markov games, see for example Littman (1994); Busoniu et al. (2008); Zhang et al. (2021). However, this line of work does not focus on the overall quality of learning outcomes. Another line of work on repeated games with such carryover effect between rounds, focusing on such overall learning outcomes, is the repeated ad-auction game with limited budgets, where the remaining budgets make the games no longer independent. See for example Balseiro and Gur (2019); Gaitonde et al. (2023); Fikioris and Tardos (2023); Fikioris et al. (2024).
2 Model
We consider a modified version of the model used by Gaitonde and Tardos (2020, 2023). There is a system of $n$ queues and $m$ servers. We divide time into discrete time steps $t=0,1,\dots$. In every step $t$, the following occurs. See the illustration in Figure 1.
  1. 1.
    Packets arrive at each queue $i$ according to a fixed probability $\lambda_{i}$. Formally, we model this by defining $B^{i}_{t}\sim\text{Bern}(\lambda_{i})$ as an independent random variable that indicates whether a packet arrives at queue $i$ at time $t$. After arrival, each queue that contains at least one packet selects a server to which it attempts to send a single packet.
  2. 2.
    Each server $j$ processes incoming packets as follows: (i) If the server’s buffer is full, all received packets are rejected and returned to their respective queues. (ii) Otherwise, the server accepts one of the received packets, chosen uniformly at random, and places it in its buffer, rejecting the remaining packets. (iii) The server then attempts to process the packet in its buffer. With probability $\mu_{j}$, the processing succeeds, and the packet is removed from the buffer. If the processing fails, the packet remains in the buffer.
  3. 3.
    Any packets not accepted into a server’s buffer are returned to their respective queues. Each queue receives bandit feedback based on the outcome of sending its packet: a reward of zero if the packet was not accepted by the server a reward of one if the packet was successfully placed in the server’s buffer. [^3] Feedback is received immediately after a packet is placed in the buffer, not necessarily when it is successfully processed.
Figure 1: A three-queue, three-server system in a single time step. Discs represent packets; left squares are queues, right squares are servers, and small squares are buffers. In this time step, two packets are sent to the top server, one of them is accepted into a buffer and then processed, the other one is returned to its queue. For the other two servers, one admits a packet into a buffer but fails to serve it and one processes a packet that was in its buffer.
Definition 1.
We say that a system with a schedule for sending packets is stable if the expected number of packets waiting to be sent or served in the system is bounded by a constant at all times. We allow the constant to depend on the number of queues or servers and the parameters of the system, but it may not grow with time $T$.
It is clear that if $\sum_{i}\lambda_{i}>\sum_{j}\mu_{j}$ then even a fully coordinated system cannot be stable, and the total number of packets waiting at the queues must grow linearly in time. In the case of a single queue with a single server, $\lambda<\mu$ is the necessary and sufficient condition for keeping the packets remaining in the system bounded by a constant (which depends on the gap in the inequality) at all times. The analogous condition is clearly necessary but not sufficient in case of multiple servers.
Lemma 1.
Consider an example with 1 queue and 2 servers, each with $\mu=\frac{1}{2}$. We claim that $\lambda<\frac{23}{24}<1$ is needed for the system to be stable even with full coordination.
Proof.
To see why this is true, consider 3 consecutive time steps. To analyze the maximum processing rate we look at the scenario where the queue sends a packet in each of these steps. We claim that with probability $1/8$ it holds that in one of the three times the queue must fail to have a packet accepted by a server. To see this, assume the following.
  • Assume that at step 1 the queue sends a packet to one of the servers, which fails to serve it, but may keep it in its buffer. This event has probability at least $\frac{1}{2}$.
  • With full coordination, the queue is aware of the first buffer being full and can try to send the next packet to the second server. Assume that at this step neither of the two servers succeed in sending packets, but as before, may accept the new packet into its buffer. This event has probability at least $\frac{1}{4}$.
  • Even if the first two packets were accepted into buffers, in this case, at the third time step, both buffers are full, and so there is no place for the queue to send.
This proves that during a set of consecutive three steps the maximum expected service rate is at most
$\frac{\frac{1}{8}\cdot 2+\frac{7}{8}\cdot 3}{3}=\frac{23}{24}.$
Thus, if the arrival rate $\lambda$ is higher, we will have a buildup of the queue linear in $T$. ∎
3 Analysis
Our main result is the following theorem:
Theorem 1.
Assuming $\sum_{i}\lambda_{i}<\frac{1}{3}\sum_{j}\mu_{j}$, $\lambda_{i}<\frac{1}{2}$ for all $i$, and all queues use a form of learning guaranteeing no-regret with high probability to identify servers they can use, the system remains stable with the expected number of packets in the system bounded by a (time-independent) constant at all times.
The key tool in our analysis is the following theorem of Pemantle and Rosenthal (1999):
Theorem 2.
Let $X_{1},X_{2},\ldots$ be a sequence of nonnegative random variables with the property that
  1. 1.
    There exist constants $\alpha,\beta>0$ such that $\mathbb{E}[X_{t+1}-X_{t}|\mathcal{F}_{t}\&X_{t}\geq\beta]<-\alpha,$ where $\mathcal{F}_{t}$ denotes the history until period $t$.
  2. 2.
    There exist $p>2$ and a constant $\theta>0$ such that for any history, $\mathbb{E}[|X_{t+1}-X_{t}|^{p}|\mathcal{F}_{t}]\leq\theta.$
Then, for any $10<r<p-$ 1, there exists an absolute constant $M=M(\alpha,\beta,\theta,p,r)$ not depending on $t$ such that $\mathbb{E}[X_{t}^{r}]\leq M$ for all $t$.
To use this theorem we will consider a long enough time in period of length $T$ and use a potential function that depends on the number of packets in each of the queues at the start of periods of length $T$. Assume that queue $i$ has $N_{i}$ packets at the start of the period. We will use the potential function $\Phi=\sum_{i}(N_{i}-(\lambda^{2}_{i}+3\delta)T)^{+}$, where $x^{+}=\max(x,0)$, and will aim to show that between the beginning and end of a $T$ long period the change in this potential satisfies the conditions of Theorem 1.
For a long enough period $T$, we expect that the packet arrivals and the servers’ service rates are all close enough to their expectations, and each queue has small enough regret in their learning algorithm. We will call a period of length $T$ goodδ for a parameter $\delta>0$ (to be chosen later) if the following three conditions are all satisfied. Later, we will also account for what happens when goodδ does not occur.
Condition 1.
The number of packets arriving to each queue $i$ in each period of length $\hat{T}\leq T$ is at most $\lambda_{i}\hat{T}+\delta T$.
Condition 2.
Each server $j$ that has a packet in its buffer (possibly one that just arrived) for at least half of the steps during this $T$-length period succeeds in serving at least $(\frac{1}{2}\mu_{j}-\delta)T$ packets.
Condition 3.
The learning algorithm of each queue $i$ accumulates regret at most $\delta T$ during the $T$-length period.
We use Chernoff and union bounds to bound from below the probability that conditions 1 and 2 hold when choosing a high enough $T$.
Lemma 2.
For any constants $\delta>0$ and $\eta>0$, by choosing the period length $T$ high enough, we can guarantee that the probability that conditions 1 or 2 fail to hold for a single queue or single server is bounded by $\eta$, and hence with probability $(1-\eta(n+m))$ the conditions hold.
To guarantee that Condition 3 holds with high probability, we will assume that queues use a learning algorithm with a high probability regret guarantee, such as EXP3.P.1, of Auer et al. (2002a), which offers a regret guarantee of $\sqrt{Tm\ln(mT/\eta)}$ with probability $1-\eta$.
Lemma 3.
For any constants $\delta>0$ and $\eta>0$, by choosing the period length $T$ high enough we can guarantee that if all queues use EXP3.P.1 as their learning algorithm, the probability that Condition 3 holds for a single queue is at least $1-\eta$. Hence, the condition holds with probability $(1-n\eta).$
Now we are ready to start proving the main result. To show the expected decrease in the potential function, we consider whether a period of length $T$ satisfied the condition of being goodδ. By Lemmas 2 and 3, we see that the probability that goodδ fails to hold is bounded by $(m+2n)\eta$. The maximum possible increase in the potential function during a period of $T$ steps is at most $nT$ (if packets arrive during each step at each of the queues, and none reach the servers), so bad events contribute at most $(m+2n)\eta nT$ to the expected change in potential over the $T$ steps.
Next, we consider the change in potential for a goodδ period of length $T$.
Lemma 4.
If all servers have a packet in their buffer (possibly one that just arrived) for more than half of the iterations, and we are in the goodδ case, the total number of packets at the queues decreases by at least $(\frac{1}{2}\sum_{j}\mu_{j}-\sum_{i}\lambda_{i}-(n+m)\delta)T$.
Proof.
By condition 2 the total number of packets served is at least $\sum_{j}(\frac{1}{2}\mu_{j}-\delta)T$. By condition 1 the total number of arriving packets is at most $\sum_{i}(\lambda_{i}+\delta)T$. Combining these two bounds establishes the lemma. ∎
Lemma 5.
Suppose there is a server whose buffer is empty more than half of the $T$ iterations, and we are in the goodδ case. Consider a queue $i$ and assume that $\delta<\frac{1}{2}(\frac{1}{2}-\lambda_{i})$, and $\delta<\lambda_{i}$. Then, the number of packets at this queue will decrease, unless the number $N^{\prime}_{i}$ left at the end the period is at most $N^{\prime}_{i}\leq(\lambda^{2}_{i}+3\delta)T$. If at the start there are at least $T$ packets in the queue, then the number decreases by at least $\left(\frac{1}{2}-(\lambda_{i}+2\delta)\right)T$.
Proof.
Consider a queue $i$. Suppose that the queue starts with $N_{i}$ packets and ends the period with $N^{\prime}_{i}$. If the server was sending a packet in every step, then by condition 3 it cleared at least $(\frac{1}{2}-\delta)T$ packets, and by condition 1 at most $(\lambda_{i}+\delta)T$ arrived, decreasing the total by at least $(\frac{1}{2}-(\lambda_{i}+2\delta))T$. This proves the second part of the claim, as $T$ packets (or more) at the start will guarantee sending a packet in every step.
Now suppose that the queue has packets in the last $\hat{T}$ steps but was empty in the previous step. Let $S_{i}$ denote the number of packets it successfully cleared in this $\hat{T}$-step period. Assume that some $(T-\hat{T})\lambda_{i}+X_{i}$ packets arrived during the first $(T-\hat{T})$ steps with $X_{i}\leq\delta T$ by Condition 1. Now by the no-regret Condition 3 we get that
$N_{i}+\lambda_{i}(T-\hat{T})+X_{i}+S_{i}\geq\frac{1}{2}(N_{i}+\lambda_{i}(T-\hat{T})+X_{i}+\hat{T})-\delta T,$
which implies (using that $X_{i}\leq\delta T$)
$N_{i}\geq\hat{T}-\lambda_{i}(T-\hat{T})-2S_{i}-3\delta T.$
By condition 1 the final number of packets in the queue at the end of the period is at most $N^{\prime}_{i}\leq\lambda_{i}\hat{T}+\delta T-S_{i}$.
Putting these two inequalities together we get that
$N_{i}-N^{\prime}_{i}$$\geq$ $\hat{T}-\lambda_{i}(T-\hat{T})-2S_{i}-3\delta T$
$-(\lambda_{i}\hat{T}+\delta T-S_{i})$
$\geq$ $\hat{T}-\lambda_{i}T-4\delta T-S_{i}.$
This shows that the number of packets at queue $i$ decreases, unless $\hat{T}\leq(\lambda_{i}+4\delta)T+S_{i}$. However, in this case we get that the number of packets at the end of the period is at most $N_{i}^{\prime}\leq\lambda_{i}\hat{T}+\delta T-S_{i}\leq(\lambda^{2}_{i}+4\lambda_{i}\delta)+\delta)T-(1-\lambda_{i})S_{i}\leq(\lambda^{2}_{i}+3\delta)T$, as $\lambda_{i}<1/2$, proving the claim. ∎
Now we are ready to prove Theorem 1.
Proof.
(Theorem 1) Recall the potential function $\Phi=\sum_{i}(N_{i}-(\lambda^{2}_{i}+3\delta)T)^{+}$. Clearly, $\Phi$ has bounded moments, as the maximum possible increase in the potential function over a period of length $T$ is when a packet arrives at every single queue during each step and no packets are accepted at any of the servers. That is a total increase of $nT$. Similarly, the maximum possible decrease in the potential function over a period of length $T$ is when no packets arrive, and each queue succeeds in sending a packet to a server at each step. That is a total decrease of $nT$.
Next, we want to prove that when $\Phi>Tn$, the potential is expected to decrease. Recall that when goodδ fails to hold the potential can increase by as much as $nT$. By Lemmas 2 and 3 this can contribute at most $\eta(m+2n)nT$ to the expectation.
To analyze the case when goodδ holds we consider two cases separately. If all servers have packets to try to serve (in their buffer, or just arriving) at least half of the time, then by Lemma 4 the total number of packets decreases by at least $(\frac{1}{2}\sum_{j}\mu_{j}-\sum_{i}\lambda_{i}-(n+m)\delta)T$. Some of this decrease may not affect the potential function, since some of the cleared packets can come from queues that contribute zero to the potential. However, this can only affect $\sum_{i}(\lambda^{2}_{i}+3\delta)T$ packets in queue $i$, and so the potential is decreasing by at least
$\Big{(}\frac{1}{2}\sum_{j}\mu_{j}-\sum_{i}(\lambda_{i}+\lambda_{i}^{2})-(4n+m)\delta\Big{)}T.$ (1)
By the assumption that $\lambda_{i}<\frac{1}{2}$ and $\sum_{j}\mu_{j}>3\sum_{i}\lambda_{i}$, this is an $\Omega(T)$ decrease, assuming we choose a small enough $\delta$.
Now consider the case where a server is open (i.e., its buffer is empty, and no new packet arrived) half the time. If $\Phi>nT$ then some queue $i$ will have at least $T$ packets. In this case, by Lemma 5 this queue will have its packet count decrease by at least $(\frac{1}{2}-\lambda_{i}-2\delta)T$, contributing this decrease to the potential function. Lemma 5 also shows that all other queues will have their contribution to the potential decrease, unless it is zero at the end of the $T$-length period. This shows a total decrease in the potential of at least $(\frac{1}{2}-\lambda_{i}-2\delta)T=\Omega(T)$, again assuming $\delta$ is small enough.
Putting together the expected increase when goodδ fails and expected decrease in the two cases when goodδ holds (and using that $\lambda_{i}\leq\frac{1}{2}$), we get that the potential function is expected to decrease after a $T$-length period by at least
$T\min\big{(}\frac{1}{2}-\lambda_{i}-2\delta,\frac{1}{2}\sum_{j}\mu_{j}-\frac{3}{2}\sum_{i}\lambda_{i}-(4n+m)\delta\big{)}$
$-T\eta(m+2n)n.$
To guarantee the required expected decrease of the potential, we choose $\delta$ small enough so both terms in the $\min$ are at least a constant (so in the case goodδ happens the potential decreases by $\Omega(T)$) and then choose $\eta$ small enough to make the total positive. ∎
Remark 1.
Note that the proof actually shows a slightly stronger statement dependent on the $\lambda_{i}$ values. Using equation (1), it is enough to require $\sum_{i}(\lambda_{i}+\lambda_{i}^{2})<\frac{1}{2}\sum_{j}\mu_{j}$ in place of $\sum_{i}\lambda_{i}<\frac{1}{3}\sum\mu_{j}$. This improves the bound when the arrival rates are small, approaching a factor of 2 in this case.
While we do not know if the factor of 3 in the main theorem is a tight worst-case bound even when arrival rates $\lambda_{i}$ are close to $\frac{1}{2}$ (in which case $\sum_{i}(\lambda_{i}+\lambda_{i}^{2})=\sum_{i}\lambda_{i}(1+\lambda_{i})=\frac{3}{2}\sum_{i}\lambda_{i}$ yields this factor), the following example shows that we need at least a factor of 2 to ensure stability.
Theorem 3.
[Lower Bound Example] Consider a queue with $1/2$ arrival rate, facing $k$ servers, each with only capacity $1/n$. Assume that the server is choosing a uniform random server at each step. This sending plan will clearly satisfy the no-regret condition of Condition 3. We claim that we need at least $n$ servers to make the solution stable proving a lower bound of 2 needed for the needed increase in server capacity.
Proof.
So with $k$ servers, the chance that when the queue sends to a server, the previous time it did so was $i$ steps ago is $1/k\cdot(1-1/k)^{i-1}$. If the previous time the queue did sent there, the chance that the server is still busy is $(1-1/n)^{i}$, so the chance that the new packet gets accepted is
$\sum_{i=1}^{\infty}1/k\cdot(1-1/k)^{i-1}(1-(1-1/n)^{i})$
$=$ $\frac{1}{k}\sum_{i=1}(1-\frac{1}{k})^{i-1}-\frac{n-1}{nk}\sum_{i=1}^{\infty}((1-\frac{1}{k})(1-\frac{1}{n}))^{i-1}$
$=$ $1-\frac{n-1}{nk}\frac{1}{1-(1-1/k)(1-1/n)}=1-\frac{n-1}{n+k-1}$
To make this stable we need this value to be above $1/2$, so we need $k>n-1$. This results in a total capacity of $\frac{k}{n}>\frac{n-1}{n}$, essentially double the arrival rate of the queue. ∎
4 Experiments
(a) Buildup for a symmetric system with 3 servers with capacity $\mu$ and 3 queues with arrival rates $\lambda$ as a function of $\lambda/\mu$.
(b) Buildup for an ensemble of 200 randomized systems with $5$ queues and $6$ servers as a function of $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$.
Figure 2: Empirical buildup of total queue sizes normalized by $n\cdot T$ as a function of the ratio $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$ across different scenarios. The dashed vertical lines mark the ratios of $1/3$ and $1/2$ from our analysis.
In the preceding section, we derived theoretical stability conditions and worst-case bounds on the buildup of queue lengths. Specifically, our proof showed that in a system with singleton buffers, whenever a queue grows large, given that a condition of the total service capacity is satisfied, a no-regret property of the server-selection process guarantees that it will shrink at a linear rate, keeping the total lengths of queues bounded. In this section, we complement these theoretical guarantees with computational experiments to observe the typical behavior of these systems beyond the worst case, and explore the empirical impact of buffers on the dynamics.
We conduct simulations using the EXP3 algorithm Auer et al. (2002b) (see also Slivkins (2019)), implemented in Python. For the simulation parameters, we use $\gamma$ for exploration rate and $T$ for the time horizon of the simulation.
4.1 Empirical Queue Buildup
(a) Probability that a randomized system with time horizon $T$ has a queue with buildup greater than $\sqrt{T}$ as a function of the ratio $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$. Shaded regions show the 95% binomial-proportion confidence interval for the probability estimation.
(b) Empirical clearing rates (the number of packets cleared by a system normalized by time horizon $T$) across $1$,$000$ samples of a system with fixed parameters. The left histogram is without buffers and the right one is with buffers.
Figure 3: A comparison of identical systems with and without buffers.
(a) Dynamics in a system with $2$ queues with $\lambda_{1}=1/3$, $\lambda_{2}=1/6$ and 3 servers with $\mu_{1}=2/3$, $\mu_{2}=2/9$, $\mu_{3}=1/9$.
(b) Probability of sending to server 1 in a system with two queues with $\lambda_{i}=1/4$, and two servers with $\mu_{i}=2/3$, compared to the Nash equilibria. Dashed lines are the time average.
(c) Total variation distance from a pure Nash equilibrium as a function of time in multiple simulations in the system from Figure 4(b). Shaded regions are 95% confidence intervals.
Figure 4: Dynamics of probability distributions and the empirical play across different scenarios.
We start by examining the relationship between the system’s arrival rate and its service capacity. The service capacity must strictly exceed the arrival rate to prevent queue buildup. This observation is seen also in typical experimental scenarios.
Figure 2 illustrates the total empirical buildup, which is the sum of all packets left in queues after $T$ iterations, normalized by $n\cdot T$ (vertical axis) as a function of the ratio $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$ (horizontal axis). Figure 2(a) focuses on a symmetric system with three servers such that $\mu_{1}=0.8$ and $\mu_{2}=\mu_{3}=0.2$ and three queues with equal arrival rates $\lambda_{i}$, which range between $0.02$ and $0.4$. The simulations run for $T=50$,$000$ steps with $\gamma=1/\sqrt{T}$. The solid line represents the average buildup over $200$ independent simulations for each value of $\lambda_{i}$, and the shaded region indicates the range between minimum and maximum buildup values observed in the simulations. A clear transition emerges: when $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}>0.9$, the buildup becomes proportional to $T$. Crucially, this transition occurs at a point above $0.5$ and below $1$ (where buildup is inevitable).
Figure 2(b) presents the same analysis for an ensemble of randomized systems with $5$ queues and $6$ servers. To generate such instances, we proceed as follows: generate $200$ base parameters by selecting 5 queue arrival rates, $\lambda_{i}$, and 6 server capacity rates, $\mu_{j}$, uniformly at random in $(0,1)$. For each value of the ratio $r=\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$ (with $r$ in $(0,1]$), rescale the arrival rates such that the ratio $\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}=r$ is satisfied. This process is repeated for each of the $200$ sets of base parameters. The simulations are run for $T=50$,$000$ and $\gamma=1/\sqrt{T}$. In Figure 2(b), the solid line represents the average buildup over the $200$ simulation for each $r$ value; the shaded region indicates the range between the $2.5$ and $97.5$ percentiles of buildup values. The results confirm that, even for asymmetric and randomized systems, the transition point remains above $0.5$ and below $1$. In both Figure 2(a) and Figure 2(b) the vertical lines indicate where $r=\frac{1}{3}$ and $r=\frac{1}{2}$. The empirical buildups in these simulations show that while the worst case bound is $\sum_{i}\lambda_{i}<\frac{1}{3}\sum_{j}\mu_{j}$, systems with learning typically do better than the worst case at clearing packets.
4.2 The Impact of Buffers
We now compare systems with vs. without buffers. The methodology follows the one described in the previous subsection. Figure 3(a) illustrates the effect of adding buffers to a system in the random-instance scenario, with $n=5$ queues and $m=6$ servers as described earlier. It shows the probability that some queue in the system has at time $T$ a buildup greater than $\sqrt{T}$ as a function of the arrival-to-capacity ratio for systems with and without buffers. The solid lines show the observed frequency in $300$ independent simulations for each value of $r=\sum_{i}\lambda_{i}/\sum_{j}\mu_{j}$ in $(0,1]$; the shaded region shows the $95\%$ confidence interval in estimating these probabilities. We observe that systems without buffers require a higher capacity relative to the arrival rate to prevent buildup.
Figure 3(b) shows the distributions of packet clearing rates for two identical systems, with and without buffers, over $1$,$000$ simulations with $T=10$,$000$ and $\gamma=1/\sqrt{T}$. There are $4$ queues with $\lambda_{1}=0.6,\lambda_{2}=\lambda_{3}=\lambda_{4}=0.3$ and $5$ servers with $\mu_{1}=0.8,\mu_{2}=\mu_{3}=0.4,\mu_{4}=\mu_{5}=0.2$. The clear separation between the distributions demonstrates that buffers increase efficiency. In the system without buffers, queues preferred the servers with higher capacity, which led to more competition and a lower clearing rate since the three high-capacity servers could not process all packets. [^4] The total service rate of the three highest-capacity servers is exactly the same as the total arrival rate, hence not enough for stability. In the system with buffers, queues do learn to extract good service rates also from the lower capacity servers, which increases the overall clearing rate of the system.
4.3 Dynamics and Convergence
Next, we look at the dynamics of the learning agents in our systems. We observe that in all parameter configurations we have tested, the dynamics converge in last iterate (up to a small noise level due to the finite time horizon and step size in our simulations) to Nash equilibria, as illustrated in Figure 4. This is surprising, as even without the carryover effects that our games have, no-regret dynamics need not converge to Nash equilibria. [^5] In fact, no-regret dynamics may fail to converge to any single coarse correlated equilibrium even in the average-iterate sense Kolumbus and Nisan (2022a, b). Figure 4(a) depicts the dynamics in systems with $2$ queues and $3$ servers. It can be seen that the probabilities of play of the algorithms approximately converge to a stationary distribution. Note that a stationary distribution of play and the no-regret condition imply that this distribution is a Nash equilibrium. Figure 4(b) shows a dynamic in a simpler system which has two pure Nash equilibria and one mixed equilibrium. As can be seen in the instance in the figure, which is a typical one, the dynamic in this system converges to a pure equilibrium. Figure 4(c) depicts the total variation distance from the closest Nash equilibrium as a function of the horizon $T$ in $200$ simulations of the same system for each $T$. The strategies quickly converge, and so the distance of the historical average of play shrinks as well, roughly as $1/\sqrt{T}$.
Acknowledgments
Éva Tardos was supported in part by AFOSR grant FA9550-23-1-0410, AFOSR grant FA9550-231-0068, and ONR MURI grant N000142412742.